OpenStudy (anonymous):

@campbell_st what does "geometric" mean Is the sequence geometric? 2, 6, 18, 54,...

OpenStudy (campbell_st):

geometric means that you multiply each term by the same number to get the next. in your sequence its 2, 2*3, 6*3, 18*3 so each term is multiplied by 3 to get the next.

OpenStudy (agent0smith):

You can find whether a sequence is geometric, by dividing a term by the term right before it. 2, 6, 18, 54 eg 6/2 = 3 18/6 = 3 54/18 = 3 so the sequence is geometric, since there is a common ratio of 3. If you divide a term by the one before it, and don't get the same ratio for all terms, then it's not geometric.

OpenStudy (anonymous):

so that would make it "yes"

OpenStudy (agent0smith):

Eg is this sequence geometric? 2, 8, 24, 48 8/2 = 4 24/8 = 3 48/24 = 2 the ratio is not the same, so it's not geometric.

OpenStudy (anonymous):

does it have to be mutilply? @campbell_st

OpenStudy (agent0smith):

Yes, it has to be the same multiple from one term to the next.

OpenStudy (agent0smith):

It can be multiplied by a fraction, though.

OpenStudy (campbell_st):

well dividing is the same as multiplying by a fraction 100, 50, 25, 12.5 is dividing by 2 or multiplying by 1/2

OpenStudy (anonymous):

okay so it can be either of the two but it cant be add or minus right?

OpenStudy (agent0smith):

Adding or subtracting the same amount would make it an arithmetic sequence.

OpenStudy (anonymous):

o okay just making sure! thank you!

OpenStudy (agent0smith):

No prob, you're welcome.